How to know the displacement produced by a force
By F=ma where a=v/t and where v = d/t
1
Expert's answer
2018-06-04T08:25:08-0400
If this is a constant force:
F=ma
a=F/m.
The displacement produced by a force is
d=(at^2)/2=F/2m t^2.